# Unfitted hybrid high-order methods stabilized by polynomial extension for elliptic interface problems

### v1

@@ Introduction
Define the discrete error as $\hat{e}_{h}:=\hat{I}_{h}^{k}(u)-\hat{u}_{h}\in\widehat{\mathcal{U}}_{h0}$. Since $a_{h}(\hat{e}_{h},\hat{e}_{h})=\Phi_{h}(\hat{e}_{h})$, invoking the stability result from Lemma 4.3 and the bound on the consistency error from Lemma 4.5 gives
$\|\hat{e}_{h}\|_{h0}^{2}\lesssim\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\epsilon_{T,i}(u_{i})^{2}.$
Invoking the triangle inequality and the approximation result from Lemma 4.2 establishes the first bound in (54). The second bound follows from the shape-regularity of the mesh, and the $H^{s}$-stability of the extension operators $E_{i}^{s}$. ∎
@@
Figure 5: Circular (top row) and flower-like (bottom row) interfaces. Left column: Coarsest mesh. Central column: Outcome of pairing procedure for polynomial extension stabilization, cells in $\mathcal{T}_{h}^{\textup{KO},1}$ are colored in green, and cells in $\mathcal{T}_{h}^{\textup{KO},2}$ are colored in blue, arrows indicate the pairing cell. Right column: Outcome of pairing procedure for cell-agglomeration stabilization, agglomerated cells are colored in dark.
In all cases, we consider the unit square domain $\Omega:=(0,1)^{2}$ discretized with uniform Cartesian meshes of size $h:=\sqrt{2}\times 0.1\times 2^{-\ell}$ with $\ell\in\{0,\ldots,4\}$. The polynomial degree is taken such that $k\in\{0,\ldots,3\}$. The interface is defined using a level-set function $\Phi$, with $\Gamma:=\{(x,y)\in\Omega\mid\Phi(x,y)=0\}$ and $\Omega_{i}=\{(x,y)\in\Omega\mid(-1)^{i}\Phi(x,y)>0\}$ for all $i\in\{1,2\}$. We always take $\kappa_{1}:=1$, and modify the diffusivity contrast by taking $\kappa_{2}=10^{m}$, $m\in\{1,\ldots,4\}$, where $m=1$ correspond to no contrast and $m=4$ to a highly contrasted setting. We consider two shapes for the interface: a circular shape and a flower-like shape, defined by the following level-set functions:
@@
Quadratures along the interface and in the cut sub-cells are realized by dividing the interface into $2^{r}$ segments, $r\in\mathbb{N}$, and creating a sub-triangulation of the two cut sub-cells. The construction of the sub-triangulation is performed as discussed in [14]; see, in particular, Figure 4.1 therein for an illustration. Unless explicitly stated otherwise, we set $r:=8$.
Another important aspect is the choice of the polynomial bases in the sub-cells and the sub-faces. Both bases are realized by considering centered and scaled monomials. In particular, the barycenter of each sub-cell is used for the centering of the corresponding basis. The barycenter is computed using the above sub-triangulation. Unless explicitly stated otherwise, the scaling of the monomials is isotropic and uses half of the diameter of the original cell.
@@
$u(x,y):=\sin(\pi x)\sin(\pi y).$ (56)
Errors are reported in Figure 6 for the circular interface (left panel) and the flower-like interface (right panel). Here and below, errors are evaluated using the energy norm considered in Theorem 4.6. In both panels, we compare the results obtained using stabilization by polynomial extension (solid lines) and by cell agglomeration (dashed lines). Both approaches yield similar results, with slightly better errors when using polynomial extension, especially on coarse meshes.
@@
A further comparison between polynomial extension and cell agglomeration is provided in Figure 8 which compares the sparsity profiles of the stiffness matrices obtained with both procedures. We observe that both approaches lead to a similar sparsity profile, an that the stiffness matrix corresponding to cell agglomeration is slightly smaller (as expected). The sparsity pattern of the matrix corresponding to cell agglomeration directly results from the data structure of the agglomerated mesh, whereas the sparsity pattern of the matrix corresponding to polynomial extension is derived at the assembly stage by means of the pairing operator.
Figure 9: Left: Errors for the exact solution (57) as a function of the diffusivity contrast on the finest mesh ($\ell=4$) and for various polynomial degrees ($k\in\{0,\ldots,3\}$). Right: Errors as a function of the mesh size for various values of the sub-triangulation parameter, $k=3$, $r\in\{2,4,6,8\}$, and various diffusivity contrasts, $m\in\{1,\ldots,4\}$.
@@
$u_{1}(\rho):=\frac{\rho^{6}}{\kappa_{1}},\qquad u_{2}(\rho):=\frac{\rho^{6}}{\kappa_{2}}+R^{6}\bigg{(}\frac{1}{\kappa_{1}}-\frac{1}{\kappa_{2}}\bigg{)}.$ (57)
@@
$u_{1}(\rho)$ $:=\frac{\rho^{6}}{\kappa_{1}},$ $\qquad u_{2}(\rho)$ $:=\frac{\rho^{8}-R^{8}}{\kappa_{2}}+\frac{R^{6}}{\kappa_{1}},$ (58a)
$u_{1}(\rho)$ $:=\frac{\rho^{6}}{\kappa_{1}},$ $\qquad u_{2}(\rho)$ $:=\frac{\rho^{6}}{\kappa_{2}}.$ (58b)
@@
Figure 11: Left: Errors as a function of the mesh size for the exact solution (59) for various polynomial degrees $k\in\{0,\ldots,3\}$ (with sub-triangulation parameter set to $r=10$). Right: Errors as a function of the mesh size for $k=3$ and $r\in\{2,4,6,8,10\}$. No diffusivity contrast.
@@
$u_{1}(x,y):=\cos(y)e^{x},\qquad u_{2}(x,y):=\sin(\pi x)\sin(\pi y),$ (59)
leading to variable jump data $g_{D}$ and $g_{N}$. Here, we do not consider a diffusivity contrast ($\kappa_{1}=\kappa_{2}=1$). Errors as a function of the mesh size are reported in the left panel of Figure 11 for various polynomial degrees $k\in\{0,\ldots,3\}$ and sub-triangulation parameter set to $r=10$. Optimal convergence rates are observed in all cases. To motivate the choice for $r$, we report in the right panel of Figure 11 the errors as a function of the mesh size for $k=3$ and $r\in\{2,4,6,8,10\}$. We observe that the geometric error does not pollute the discretization error only for $r=10$ (notice in passing the very low values attained by the error).
@@
Finally, we perform a brief study of the conditioning of the stiffness matrix. In the left panel of Figure 12, we consider a circular interface with radius $R=\frac{1}{3}+\frac{i}{32}$, $i\in\{-4,\ldots,4\}$, and we report the Euclidean condition number as a function of the radius, using either polynomial extension (solid lines) or cell agglomeration (dashed lines). We consider the coarsest mesh ($\ell=0$) and polynomial degrees $k\in\{0,\ldots,3\}$. We observe that, for each polynomial degree, the conditioning remains fairly insensitive to the stabilization method of the ill-cut cells. To exacerbate the effect of ill-cut cells, we now consider a square interface with level-set function $\Phi_{\textsc{s}}(x,y):=\max(x-a,y-b)-(0.25+\delta)$ and position parameter $\delta=0.5\times 10^{-p}$, $p\in\{1,\ldots,5\}$. Since cut cells can become very anisotropic, we use here an anisotropic scaling of the polynomial bases in each sub-cell, whereby the variable $x$ (resp., $y$) is scaled by $\frac{1}{2}h_{x}$ (resp., $\frac{1}{2}h_{y}$), where $(h_{x},h_{y})$ are the lengths of the sides of the bounding box of the sub-cell. The right panel of Figure 12 reports the Euclidean condition number as a function of the position parameter of the square interface. We observe that the use of polynomial extension leads to some sensitivity of the conditioning for severe ill cuts, with a linear growth in $\delta^{-1}$. Similar observations are reported in [12] in the context of isogeometric methods. Somewhat surprisingly, the cell agglomeration approach appears to deliver robust condition numbers. We emphasize that the errors still behave optimally when using polynomial extension (figure omitted for brevity), in agreement with the above convergence analysis.
@@
(1) Proof of (38a). Let $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$ and $\phi\in\mathbb{P}^{\ell}(T^{i};\mathbb{R})$. Invoking the discrete trace inequality from [18, Lemma 3.4] shows that
@@
where the last bound follows from the construction of the pairing operator. We next observe that there is a ball $\mathcal{B}(\Delta(T))$ of diameter $\gamma h_{T}$ such that $\Delta(T)\subset\mathcal{B}(\Delta(T))$, where $\gamma$ only depends on the mesh shape-regularity parameter. We also recall that, since $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$, there is a ball $\mathcal{B}(T,i)$ of diameter $\vartheta h_{T}$, so that $\mathcal{B}(T,i)\subset T^{i}$. We then infer that
@@
where the inverse inequality invoked in the second inequality follows by the arguments given in the proof of [14, Lemma 3.4]. This completes the proof of (38a).
@@
where we also used the mesh shape-regularity which implies that $h_{T}\lesssim h_{S}\lesssim h_{T}$ for all $S\in\Delta(T)$. The bound (38b) now follows from the discrete Poincaré inequality established in [14, Lemma 3.4]. ∎
@@
We have $\|\tilde{v}_{i}-\langle v_{i}\rangle_{\mathcal{B}(T,i)}\|_{S}\leq\|\tilde{v}_{i}-\langle v_{i}\rangle_{\mathcal{B}(T,i)}\|_{\Delta(T)}$ since, by construction, $S\subset\Delta(T)$. Let $\langle\tilde{v}_{i}\rangle_{\Delta(T)}$ denote the mean-value of $\tilde{v}_{i}$ on $\Delta(T)$. Invoking [27, Lemma 5.7] gives
@@
(2) The higher-order version of (60) is established by invoking the Morrey polynomial of $v_{i}$ based on mean-values of higher-order derivatives of $v_{i}$ (see, e.g., the proof of [27, Lemma 5.6]). Omitting the details for brevity, this gives, for all $0\leq m<s$, a polynomial $q_{m}(v_{i})\in\mathbb{P}^{k+1}(T^{i};\mathbb{R})$ such that, for all $S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)$,
@@
(4) Invoking a multiplicative trace inequality on all the faces composing $(\partial S)^{i}$ for all $S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)$, and since $s>\frac{3}{2}$ by assumption, we infer from the above bounds that
@@
(5) To prove (41b), we bound the jump across $S^{\Gamma}$ by the triangle inequality, so that we need to estimate the traces on $S^{\Gamma}$ from both sides of $S^{\Gamma}$, for all $S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)$. On each side, we invoke a multiplicative trace inequality on $S^{\Gamma}$. This inequality is established with a slight adaptation of the arguments in the proof of [18, Lemma 3.3], whereby we make a specific choice for the apex of the cone, say $C(S^{\Gamma})$, considered in that proof.

### v2

@@ Introduction
Define the discrete error as $\hat{e}_{h}:=\hat{I}_{h}^{k}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}})-\hat{u}_{h}\in\widehat{\mathcal{U}}_{h0}$. Since $a_{h}(\hat{e}_{h},\hat{e}_{h})=\Phi_{h}(\hat{e}_{h})$, invoking the stability result from Lemma 4.4 and the bound on the consistency error from Lemma 4.6 gives
$\|\hat{e}_{h}\|_{h0}^{2}\lesssim\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\big\{\epsilon_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(u^{\textup{ex}})}\big\}^{2}.$
Owing to the approximation result from Lemma 4.3 and since $\kappa_{1}\leq\kappa_{2}$ and $\Delta(T)\subset\Delta_{2}(T)$, we infer that
$\|\hat{e}_{h}\|_{h0}^{2}\lesssim\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{2(s-1)}|E_{i}^{s}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})|_{H^{s}(\Delta_{2}(T))}^{2}.$
Invoking the triangle inequality then establishes the first bound in (54). The second bound follows from the shape-regularity of the mesh, and the $H^{s}$-stability of the extension operators $E_{i}^{s}$. ∎
@@
Figure 5: Circular (top row) and flower-like (bottom row) interfaces. Left column: Coarsest mesh. Central column: Outcome of pairing procedure for polynomial extension stabilization, cells in $\mathcal{T}_{h}^{\textup{KO},1}$ are colored in green, and cells in $\mathcal{T}_{h}^{\textup{KO},2}$ are colored in blue, arrows indicate the pairing operator. Right column: Outcome of pairing procedure for cell-agglomeration stabilization, agglomerated cells are colored in dark.
In all cases, we consider the unit square domain $\Omega:=(0,1)^{2}$ discretized with uniform Cartesian meshes of size $h:=\sqrt{2}\times 0.1\times 2^{-\ell}$ with $\ell\in\{0,\ldots,4\}$. The polynomial degree is taken such that $k\in\{0,\ldots,3\}$. The interface is defined using a level-set function $\Phi$, with $\Gamma:=\{(x,y)\in\Omega\mid\Phi(x,y)=0\}$ and $\Omega_{i}=\{(x,y)\in\Omega\mid(-1)^{i}\Phi(x,y)>0\}$ for all $i\in\{1,2\}$. We always take $\kappa_{1}:=1$, and modify the diffusivity contrast by taking $\kappa_{2}=10^{m}$, $m\in\{0,\ldots,4\}$, where $m=0$ corresponds to no contrast and $m=4$ to a highly contrasted setting. We consider two shapes for the interface: a circular shape and a flower-like shape, defined by the following level-set functions:
@@
Quadratures along the interface and in the cut sub-cells are realized by dividing the interface into $2^{r}$ segments, $r\in\mathbb{N}$, and creating a sub-triangulation of the two cut sub-cells. The construction of the sub-triangulation is performed as discussed in [12]; see, in particular, Figure 4.1 therein for an illustration. Unless explicitly stated otherwise, we set $r:=8$. We also take $\eta_{\mathcal{N}}:=20$ for the stabilization bilinear form $s_{h}^{\mathcal{N}}$ in (30). Concerning the flagging of ill-cut cells, we use the parameter $\vartheta:=0.3$. Our numerical experiments below show a marginal impact of the choice of this parameter on stability and accuracy. As this numerical observation is not backed up by theoretical results, we recommend to use a flagging parameter $\vartheta$ as indicated above.
Another important aspect is the choice of the polynomial bases in the sub-cells and the sub-faces. Both bases are realized by considering centered and scaled monomials. In particular, the barycenter of each sub-cell is used for the centering of the corresponding basis. The barycenter is computed using the above sub-triangulation. The scaling of the monomials is isotropic and uses half of the diameter of the original cell. In the ill-cut cells, say $S_{i}$ with $i\in\{1,2\}$, the centering and scaling is defined by considering the merged cell $S_{i}\cup\mathcal{N}_{i}(S_{i})$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}(x,y):=\sin(\pi x)\sin(\pi y).$ (56)
Errors are reported in Figure 6 for the circular interface (left panel) and the flower-like interface (right panel). Here and below, errors are evaluated using the energy norm considered in Theorem 4.7. In both panels, we compare the results obtained using stabilization by polynomial extension (solid lines) and by cell agglomeration (dashed lines). Both approaches yield similar results, with slightly better errors when using polynomial extension, especially on coarse meshes, except for $k=3$ and the flower-like interface.
@@
A further comparison between polynomial extension and cell agglomeration is provided in Figure 8 which compares the sparsity profiles of the stiffness matrices obtained with both procedures using the lowest-order polynomial setting. We observe that both approaches lead to a similar sparsity profile, and that the stiffness matrix corresponding to cell agglomeration is slightly smaller (as expected). The sparsity pattern of the matrix corresponding to cell agglomeration directly results from the data structure of the agglomerated mesh, whereas the sparsity pattern of the matrix corresponding to polynomial extension is derived at the assembly stage by means of the pairing operator.
Figure 9: Left: Errors for the exact solution (57) as a function of the diffusivity contrast on the finest mesh ($\ell=4$) and for various polynomial degrees ($k\in\{0,\ldots,3\}$). Right: Errors as a function of the mesh size for various values of the sub-triangulation parameter, $k=3$, $r\in\{2,4,6,8\}$, and various diffusivity contrasts, $m\in\{0,\ldots,4\}$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}(\rho):=\frac{\rho^{6}}{\kappa_{1}},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{2}}(\rho):=\frac{\rho^{6}}{\kappa_{2}}+R^{6}\bigg(\frac{1}{\kappa_{1}}-\frac{1}{\kappa_{2}}\bigg).$ (57)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}(\rho)$ $:=\frac{\rho^{6}}{\kappa_{1}},$ $\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{2}}(\rho)$ $:=\frac{\rho^{8}-R^{8}}{\kappa_{2}}+\frac{R^{6}}{\kappa_{1}},$ (58a)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}(\rho)$ $:=\frac{\rho^{6}}{\kappa_{1}},$ $\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{2}}(\rho)$ $:=\frac{\rho^{6}}{\kappa_{2}}.$ (58b)
@@
Figure 11: Left: Errors as a function of the mesh size for the exact solution (59) for various polynomial degrees $k\in\{0,\ldots,3\}$ (with sub-triangulation parameter set to $r=10$). Right: Errors as a function of the mesh size for $k=3$ and $r\in\{4,6,8,10,11\}$. No diffusivity contrast.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}(x,y):=\cos(y)e^{x},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{2}}(x,y):=\sin(\pi x)\sin(\pi y),$ (59)
leading to variable jump data $g_{D}$ and $g_{N}$. Here, we do not consider a diffusivity contrast ($\kappa_{1}=\kappa_{2}=1$). Errors as a function of the mesh size are reported in the left panel of Figure 11 for various polynomial degrees $k\in\{0,\ldots,3\}$ and sub-triangulation parameter set to $r=10$. Optimal convergence rates are observed in all cases. To motivate the choice for $r$, we report in the right panel of Figure 11 the errors as a function of the mesh size for $k=3$ and $r\in\{4,6,8,10,11\}$. We observe that the geometric error does not pollute the discretization error only for $r=10$, except for $k=3$ on the finest mesh where $r=11$ is actually necessary to suppress the geometric error (notice in passing the very low values attained by the error).
@@
Finally, we perform a brief study of the conditioning of the stiffness matrix. In the left panel of Figure 12, we consider a circular interface with radius $R:=\frac{1}{3}+\frac{i}{32}$, $i\in\{-4,\ldots,4\}$, and we report the Euclidean condition number as a function of the radius, using either polynomial extension (solid lines) or cell agglomeration (dashed lines). We consider the coarsest mesh ($\ell=0$) and polynomial degrees $k\in\{0,\ldots,3\}$. We observe that, for each polynomial degree, the conditioning remains fairly insensitive to the stabilization method of the ill-cut cells. To exacerbate the effect of ill-cut cells, we now consider a square interface with level-set function $\Phi_{\textsc{s}}(x,y):=\max(x-a,y-b)-(0.25+\delta)$ and position parameter $\delta:=0.5\times 10^{-p}$, $p\in\{2,\ldots,9\}$. The right panel of Figure 12 reports the Euclidean condition number as a function of the position parameter of the square interface. We observe full robustness of the condition number with respect to the cut position, for both polynomial extension and cell agglomeration. Further numerical results (not displayed for brevity) indicate that a variant of the stabilization bilinear form $s_{h}^{\mathcal{N}}$ where the penalty is enforced on the faces of $T$ with a weight $h_{T}^{-1}$ does not lead to a robust behavior, but instead to a linear growth in $\delta^{-1}$ for $k\in\{2,3\}$. We also emphasize that the errors still behave optimally (figure omitted for brevity), in agreement with the above convergence analysis.
@@
(1) Proof of (38a). Let $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$ and $\phi\in\mathbb{P}^{\ell}(T^{i};\mathbb{R})$. We have
@@
where the first bound follows from the discrete trace inequality from [16, Lemma 3.4] upon observing that the aggregated cell $T\cup\mathcal{N}_{i}^{-1}(T)$ satisfies the ball condition invoked therein, and the second bound follows from the construction of the pairing operator. We next observe that there is a ball $\mathcal{B}(\Delta(T))$ of diameter $\gamma h_{T}$ such that $\Delta(T)\subset\mathcal{B}(\Delta(T))$, where $\gamma$ only depends on the mesh shape-regularity parameter. We also recall that, since $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$, there is a ball $\mathcal{B}(T,i)$ of diameter $\vartheta h_{T}$, so that $\mathcal{B}(T,i)\subset T^{i}$. We then infer that
@@
where the inverse inequality invoked in the second inequality follows by the arguments given in the proof of [12, Lemma 3.4]. This completes the proof of (38a).
@@
where we also used the mesh shape-regularity which implies that $h_{T}\lesssim h_{S}\lesssim h_{T}$ for all $S\in\Delta(T)$. The bound (38b) now follows from the discrete Poincaré inequality established in [12, Lemma 3.4]. ∎
@@
We have $\|\tilde{v}_{i}-\langle v_{i}\rangle_{\mathcal{B}(T,i)}\|_{S}\leq\|\tilde{v}_{i}-\langle v_{i}\rangle_{\mathcal{B}(T,i)}\|_{\Delta(T)}$ since, by construction, $S\subset\Delta(T)$. Let $\langle\tilde{v}_{i}\rangle_{\Delta(T)}$ denote the mean-value of $\tilde{v}_{i}$ on $\Delta(T)$. Invoking [24, Lemma 5.7] gives
@@
(2) The higher-order version of (60) is established by invoking the Morrey polynomial of $v_{i}$ based on mean-values of higher-order derivatives of $v_{i}$ (see, e.g., the proof of [24, Lemma 5.6]). Omitting the details for brevity, this gives, for all $0\leq m<s$, a polynomial $q_{m}(v_{i})\in\mathbb{P}^{k+1}(T^{i};\mathbb{R})$ such that, for all $S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)$,
@@
(4) Invoking a multiplicative trace inequality on all the faces composing $(\partial S)^{i}$ for all $S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)$ (observe that the aggregated cell $T\cup\mathcal{N}_{i}^{-1}(T)$ satisfies the required properties), and since $s>\frac{3}{2}$ by assumption, we infer from the above bounds that
@@
(5) To prove (41b), we bound the jump across $S^{\Gamma}$ by the triangle inequality, so that we need to estimate the traces on $S^{\Gamma}$ from both sides of $S^{\Gamma}$, for all $S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)$. On each side, we invoke a multiplicative trace inequality on $S^{\Gamma}$. This inequality is established with a slight adaptation of the arguments in the proof of [16, Lemma 3.3], whereby we make a specific choice for the apex of the cone, say $C(S^{\Gamma})$, considered in that proof.